write and equation of the line passing through (0,2) that is perpendicular to y= -4x+6
Answer by rfadrogane(214)

write and equation of the line passing through (0,2) that is perpendicular to y= -4x+6.
Sol'n:
y = -4x+6
slope of the line (m).
m1 = -4
slope the line perpendicular to y = -4x+6.
m2 = 1/4, because for two perpendicular line the product of their slope is -1.
so, when this line pass through point (0,2).
Using point-slope intercept form.
y-y1=m(x-x1)
y-2 = (1/4)(x-0)
4(y-2)=x
4y-8=x
x-4y+8=0 ---answer